I need your help please and I certainly do thank you.
Solve for x to Four decimal places
18.)10^x=153
19.)1.02^4x=2

Answer by ankor@dixie-net.com(12692)

18.) 10^x = 153
Use the log equiv of exponents. Find the log of 153, we know the log of 10 is 1
x*(1) = 2.1847
:
Check on a good calc, enter: 10^2.1847 = 153.0030
:
:
The same procedure with this one
19.) 1.02^4x=2
4x * log(1.02) = log(2)
4x * .0086 = .301
.0344x = .301
x = .301/.0344
x = 8.75
:
Check on a good calc; enter: 1.02^(4*8.75) = 1.9999 ~ 2
:
Did this help?